# Evolution of reconnection flux during eruption of magnetic flux ropes

### v2

@@ 4 Results
Our simulation produces two homologous CME eruptions; each of the eruption characterized by a sudden rise in the kinetic energy and corresponding release of the magnetic energy. We first briefly discussed the initiation and evolution of the eruptions and then analyze the velocity and reconnection flux of different CME eruptions in the simulation. Note that the magnetic reconnection in our model using MHD approximation is of numerical origin, we cannot resolve length scales smaller than the Larmor radius of the proton or the mean free path. However, our goal here is to find a relation (if any exists) between the reconnection flux and the speed of the ejecta in our simulations.
4.1 Overview of eruptions
@@
Figure 9: Identification of current sheet and field lines of the flux rope from different viewing angles. The current sheet (in white) traced from the temperature isosurface with a value $\log_{10}T=6.6$ shown in panel (a). The field lines passing through such sheets gives the reconnection flux. The sigmoid fieldlines are shown in panels (b) and (c). The time in all these panels are taken at $t=27.32$ hours.
@@
The simulation starts with a magnetic flux rope (MFR) emerging from the lower boundary and moving into an existing coronal potential arcade field. The normal magnetic flux distributions at the lower boundary are represented by bipolar bands $B_{r}(R_{\odot},\theta,\phi)=B_{s}(\theta)$ where the $B_{s}(\theta)$ denotes the potential arcade field at the lower boundary, as shown in Figure (7). The emergence process is halted once a specified amount of twisted flux is driven into the corona. Although the imposed emergence of a twisted flux rope through the lower boundary may not perfectly reflect real-world conditions, it serves as a means to achieve a sequence of near-force-free coronal flux rope equilibria with increasing amounts of locally detached, twisted flux. In all cases, we observe the development of current layers with a sigmoid morphology beneath the flux rope. Magnetic reconnection within these current layers continues to add flux to the flux rope, even though the total magnetic energy gradually decreases due to reconnections. This continuous addition of flux facilitates successive eruptions.
@@
Figure 10: Time evolution of velocity, (a)-(b), and reconnection flux, (c)-(d), for both the eruptions in the simulation. The velocity is calculated by tracking the flux rope in the central meridional plane. The reconnection flux is calculated by the magnetic flux passing through the footpoints of the fieldlines as well as through the current sheet shown in Fig. 9a. The red (blue) circles are used to represent simulation data before (after) the eruption.
The first eruption is triggered when the twist of the winding of the flux rope field lines about the torus axis reaches approximately 1 wind between the anchored foot points. We stopped the emergence of the twisted torus tube at 33.5 hours when the extent of the flux rope that emerged is such that the field line’s winding about the flux rope’s axis is 1.5 between anchored footpoints. This is estimated in a simple way by multiplying the total number of winds around the circular torus, i.e., 5 winds, by the fraction of the torus circumference above the solar surface. The second eruption began after we stopped the flux emergence. The helical kink instability is expected to develop in a line-tied coronal flux rope if the total windings of the field line twist around the axis exceed a critical value between the line-tied ends (Hood & Priest, 1981; Török & Kliem, 2003; Török et al., 2004). This critical value is 1.25 (Hood & Priest, 1981) for uniformly cylindrical flux tube. The animation (corresponding to Figure 7) clearly shows significant writhing motion at the start of the eruption for the second eruption but not in first eruption, indicative of the helical kink instability. Figure 8 illustrates the evolution of kinetic and magnetic energies for both eruptions. The time at which the kinetic energy reaches its peak is designated as the eruption time.
Furthermore, we examined whether the flux rope is also unstable to the torus instability (Bateman, 1978; Kliem & Török, 2006; Isenberg & Forbes, 2007), which is triggered by a vertical force imbalance between the upward hoop force and the downward strapping force acting on the magnetic flux rope (MFR). The external strapping field is assumed to be the potential field $B_{p}$ with the same normal distribution at the lower boundary. The rate of decline with height is measured by the decay index $n=-d\ln B_{h}/d\ln h$. The critical decay index for a 3D line-tied arched flux rope has been calculated by Isenberg & Forbes (2007) to be close to 1.7. In this work, for a decay index of 1.7, the critical heights are 1.21 $R_{\odot}$ and 1.28 $R_{\odot}$, respectively for the two successive eruptions. The height of the apex of the axial field line of the coronal flux rope, estimating it to be 1.6 $R_{\odot}$ and 1.7 $R_{\odot}$ for the two successive CMEs at the time of eruption calculated by tracking the flux rope over the course of their evolution (see Appendix). For both eruptions, the flux rope crosses the critical threshold of height, indicating that it is also torus unstable. This indicates the fundamental role of kink instability and torus instability in triggering ejective eruptions of a 3D line-tied coronal flux rope.
4.2 Evolution of velocity and reconnection flux
@@
The panels (a) and (b) of Figure 10 illustrate the radial velocity of the flux rope during both eruptions. These velocities were determined by tracking the apex of each flux rope throughout their evolution (see Appendix). The first ejection has a peak radial velocity of $224\,\mathrm{km\,s^{-1}}$, which is slightly higher than the $213\,\mathrm{km\,s^{-1}}$ velocity of the second CME. This observation is understood by the time versus kinetic energy plot, which shows the second CME registered lower kinetic energy than its precursor. The initial eruption occurs at a radial distance of $1.9R_{\odot}$, while the second eruption takes place at $1.8R_{\odot}$. Due to the lower velocity of the second eruption, there is no interaction between the two CMEs. An interesting observation is that the velocity starts to decrease after the eruption which is also observed in observational data. This could be attributed to the absence of solar wind, which normally drives the magnetic flux rope (MFR) further into the heliosphere post-eruption.
The panels (c) and (d) of Figure 10 depicts the changes in reconnection flux over time for both eruptions. We monitored the hottest region beneath the flux rope in the $r\textrm{-}\theta$ plane both before and after the eruption as shown in Figure 9). This region corresponds to the current sheet, as the current sheet typically exhibits higher temperatures compared to the surrounding areas. By tracking the same hottest region, we can effectively identify the current sheet and quantify the flux passing through it. The total reconnection flux passing through the is given by
[…]
@@
where $B_{n}$ represents the magnetic flux normal to the elemental area $dS$ of the current sheet, and $B_{r}$ is the radial magnetic field in the ribbons, which are the footpoints of newly reconnected magnetic field lines at the lower boundary. The term $dA$ denotes the elemental area of these footpoints. Thus reconnection flux is determined by calculating the magnetic flux swept by the field lines passing the lower boundary. Direct measurement of $B_{n}$ and $dS$ within the current sheet is not feasible, even in simulations, due to the intricate structure of the reconnection sheet. However, $B_{r}$ and $dA$ are relatively straightforward to calculate, even for a spherical wedge-shaped domain. Therefore, this approach provides an indirect but well-defined measure of the magnetic flux passing through the reconnection sheet. This method allows for an effective evaluation of the magnetic reconnection process by focusing on the measurable quantities at the footpoints, offering valuable insights into the dynamics of the reconnection events.
@@
Figure 11: (a) CME reconnection flux vs velocity plot for the first eruption. The red (blue) circles are used to represent before (after) the eruption, respectively. The solid (dashed) line is the linear fit to the points before (after) the eruption, respectively. (b) Same as (a) but for the second eruption.
The peak reconnection flux (RC flux) for the first and second eruptions are $9\times 10^{20}$ Mx and $8.5\times 10^{20}$ Mx, respectively. The panels (c) and (d) of Figure 10 indicate that the reconnection flux decreases after reaching its peak value, with the eruptions occurring near the maximum reconnection flux. This phenomenon might be explained by the intricacies of calculating the RC flux. When an eruption occurs, the RC sheet breaks down, making it challenging to track the field lines passing through the current sheet (CS). Despite this complexity, similar patterns are observed in observational data (see panel (b) of Figure 4, suggesting a consistent underlying process. The observed decrease in reconnection flux post-peak might be tied to the dynamic changes within the RC sheet during an eruption. As the RC sheet disintegrates, accurately measuring and tracking the magnetic field lines through the CS becomes challenging. This complexity impacts the calculated values of RC flux, reflecting the real-time physical changes occurring during an eruption. Moreover, these findings align with observational data, further validating the temporal relation between the reconnection flux and the onset of eruptions (panels b-d of Fig. 4). This highlights the importance of considering both the peak and the subsequent decrease in reconnection flux when studying the dynamics of solar eruptions and magnetic reconnection events.
4.3 Correlation between MFR velocity and reconnection flux
@@
The panels (a) and (b) of Figure 11 show scatter plots of the magnetic flux rope (MFR) radial velocity ($v_{r}$) versus the reconnection flux ($\phi_{RC}$). Both before and after the eruption, the radial velocity exhibits a linear relationship with the reconnection flux for the eruptive events. Notably, the Pearson’s correlation coefficients before the eruption for the two events are 0.86 and 0.96, respectively, indicating a strong correlation between the reconnection flux and the radial velocity. This strong linear relationship persists post-eruption, with the Pearson’s correlation coefficients slightly decreasing to 0.82 and 0.91, respectively. This indicates the significant influence of reconnection flux on the radial velocity of CMEs. A similar qualitative relationship is also observed in the case of observational data (see panel (d) of Figure 4) where the velocity of the erupting filament and the associated reconnection flux simultaneously increase as indicated by the cyan shaded region. Due to observational constraints (limited time-cadence of STEREO EUVI), it was challenging to gather more data points during the initial acceleration phase of the eruption. After the peak acceleration phase, both the velocity and reconnection flux exhibit a monotonically decreasing temporal profile while maintaining similar correlated behavior. The numerical analysis from our simulation complements the above-mentioned observational results, allowing us to determine the temporal evolution of the reconnection flux and velocity and their correlation.
5 Summary and Discussion
@@
Through a simple and efficient MHD simulation, we demonstrated that homologous CMEs can result from the repeated formation and partial eruption of unstable flux ropes during the sustained emergence of highly twisted magnetic fields in an active region. Our results show that, with the continuous emergence of flux, a new current sheet (CS) forms above the same polarity inversion line (PIL) after the completion of a previous eruption, leading to a new eruption. This recurring formation and disruption of current sheet, driven by the ongoing emergence of flux within a single bipolar configuration, results in homologous eruptions.
@@
We initiated our simulation with a twisted magnetic flux rope emerging from the lower boundary into a pre-existing coronal arcade structure. Our results revealed that the flux rope initially enters a phase of slow, quasi-static ascent. During this phase, we observed the formation of a sigmoid-shaped current layer beneath the flux rope. This current layer is significant as it likely serves as the location where thin current sheets develop and magnetic reconnection occurs (Aulanier et al., 2005; Titov, 2007; Savcheva et al., 2012). The reconnection in the current layer effectively adds twisted flux to the magnetic flux rope, allowing it to rise quasi-statically before undergoing a partial eruption. Subsequent flux emergence forms a new current channel, which exhibits similar behavior and erupts after the emergence stops.
In both eruptions, the magnetic flux rope (MFR) demonstrates torus instability after reaching a height $\sim 1.6R_{\odot}$. The first eruption remains stable to helical kink instability, while the second eruption is kink unstable. The subsequent CME, in our simulation, has a slighly lower speed and lower kinetic energy. Initially, the reconnection flux increases, followed by a gradual decrease, after each eruption. These trends are consistent with the observation of an eruptive event. Our simulation indicates that the reconnection process significantly influences the evolution of the MFR. The reconnection flux shows a linear correlation with the velocity, both before and after the eruption. This result indicates that reconnection may play an important role in propelling the magnetic ejecta after the loss of equilibrium phase.
@@
Although our numerical simulation model does not include the solar wind, the inclusion of radiative cooling and explicit coronal heating makes it more realistic than previously published homologous solar eruption models (Chatterjee & Fan, 2013; Bian et al., 2022). We account for localized heating due to the formation and dissipation of current sheets in the corona, as well as the redistribution of heat through field-aligned conduction. The lower boundary of our simulation domain is positioned at the coronal base rather than the chromosphere, and therefore, our model does not simulate prominence formation. In our computational models, we incorporated an adiabatic index of $\gamma=5/3$, leads to more pronounced adiabatic cooling as the flux rope expands. Additionally, we explicitly accounted for heating processes and heat transport mechanisms. Despite these more stringent conditions, our results show that the erupting flux rope can still attain velocities characteristic of those observed in the lower corona.
@@
Although we employ an analytical configuration of the magnetic flux rope (MFR) and include various physical processes in our model, this approach may not fully represent the realistic conditions of the solar eruptions. To address the complexity and evolution of the magnetic configuration in a real scenario, a detailed and accurate description of the evolving magnetic field is essential. Our future goal is to conducting detailed analysis of solar eruptions that are directly driven or constrained by photospheric magnetograms is crucial for obtaining an accurate understanding.

### v3

@@ III Results
Our simulation produces two homologous CME eruptions; each of the eruption characterized by an impulsive increase of the kinetic energy and corresponding release of the magnetic energy. We first briefly discussed the initiation and evolution of the eruptions and then analyze the velocity, acceleration and reconnection flux of different CME eruptions in the simulation. Note that the magnetic reconnection in our model using MHD approximation is of numerical origin, we cannot resolve length scales smaller than the Larmor radius of the proton or the mean free path. However, our goal here is to find if any relation exists between the rate of change of reconnection flux and the acceleration of the ejecta in our simulations.
III.1 Synthetic eruptions
@@
Figure 5: Identification of current sheet and field lines of the flux rope from different viewing angles. The current sheet (in white) traced from the temperature isosurface with a value $\log_{10}T=6.6$ shown in panel (a). The height of the current sheet iso-surface (white) from lower boundary is about $1.4R_{\odot}$. The field lines passing through such sheets gives the reconnection flux. The sigmoid fieldlines are shown in panels (b) and (c). The time in all these panels are taken at $t=27.32$ hours.
@@
The simulation begins with a magnetic flux rope (MFR) emerging from the lower boundary and pushing into a pre-existing coronal potential arcade field. The normal magnetic flux distributions at the lower boundary are represented by bipolar bands $B_{r}(R_{\odot},\theta,\phi)=B_{s}(\theta)$ where the $B_{s}(\theta)$ denotes the potential arcade field at the lower boundary, as shown in Figure (3). The emergence process is halted once a specified amount of twisted flux is driven into the corona. Although the imposed emergence of a twisted flux rope through the lower boundary may not perfectly reflect real-sun conditions, it serves as a means to achieve a sequence of near-force-free coronal flux rope equilibria with increasing amounts of locally detached, twisted flux. In all cases, we observe the development of current layers with a sigmoid morphology beneath the flux rope. Magnetic reconnection within these current layers continues to add flux to the flux rope, even though the total magnetic energy gradually decreases due to reconnections. This continuous addition of flux facilitates successive eruptions.
@@
Figure 6: (a) Height time profile of the core of the flux rope before and after the first eruption at $t=27.9$ hours. The height is calculated by tracking the dark region (flux rope) in the central meridional plane. (b) Same as (a), but for the second eruption at $t=34.7$ hours. Time evolution of velocity, (c)-(d), and acceleration, (e)-(f), for both the eruptions in the simulation. The red (blue) circles are used to represent simulation data before (after) the eruption.
The first and second eruptions are triggered at $t=$ 27.7 hr and 34.8 hr. Figure 4 illustrates the evolution of kinetic and magnetic energies for both eruptions. The time at which the kinetic energy reaches its peak is designated as the eruption time.
@@
The panels (c) and (d) of Figure 6 illustrate the radial velocity of the flux rope during both eruptions. These velocities were determined by tracking the flux rope axis throughout their evolution. The height-time plot is produced by tracking the dark region in the $r-\theta$ plane of the temperature profile, which represents the MFR. By following this dark region consistently, we determine the MFR’s velocity during the flux rope’s evolution. The first ejection has a peak radial velocity of $224\,\mathrm{km\,s^{-1}}$, which is slightly higher than the $213\,\mathrm{km\,s^{-1}}$ velocity of the second CME. This observation is understood by the time versus kinetic energy plot, which shows the second CME registered lower kinetic energy than its precursor. The initial eruption occurs at a radial distance of $1.9R_{\odot}$, while the second eruption takes place at $1.8R_{\odot}$. Due to the lower velocity of the second eruption, there is no interaction between the two CMEs. An interesting observation is that the velocity starts to decrease after the eruption.
The reconnection flux (RC flux) passing through the current sheet is given by
$\phi_{RC}=\int B_{n}\,dS=\int B_{r}\,dA$
@@
where $B_{n}$ represents the magnetic flux normal to the elemental area $dS$ of the current sheet, and $B_{r}$ is the radial magnetic field in the ribbons, which are the footpoints of newly reconnected magnetic field lines near the lower boundary shown in Fig. 5. The term $dA$ denotes the elemental area of these footprints. Thus reconnection flux is determined by calculating the magnetic flux swept by the field lines passing near the lower boundary. Direct measurement of $B_{n}$ and $dS$ within the current sheet is not feasible, even in simulations, due to the intricate structure of the reconnection sheet. However, $B_{r}$ and $dA$ are relatively straightforward to calculate, even for a spherical wedge-shaped domain. Therefore, this approach provides an indirect but well-defined measure of the magnetic flux passing through the reconnection sheet. This method allows for an effective evaluation of the magnetic reconnection process by focusing on the measurable quantities at the footpoints, offering valuable insights into the dynamics of the reconnection events.
@@
Figure 7: Time evolution of reconnection rate (a)-(b). The reconnection flux is calculated by the magnetic flux passing through the footpoints of the fieldlines as well as through the current sheet shown in Fig. 5a. The red (blue) circles are used to represent simulation data before (after) the eruption. The smoothing is done by cubic spline. (c) CME reconnection rate vs acceleration plot for the first eruption. The red (blue) line is the linear fit to the points before (after) the eruption, respectively. (d) Same as (c) but for the second eruption.
Figure 8: Different phases of the CME eruption on 2011 August 04 as observed by STEREO-A. (a) Base difference image of the Sun in STEREO-A EUVI 171 Å . Location of STEREO spacecraft and Earth on 2011 August 4. shown in inset. (b) COR1 base-difference image superimposed with STEREO-A EUVI 171 Å (̇c) Superimposed base-difference images of STEREO-A EUVI 171 Å , COR1 and COR2. The red arrows in panels (a) and (b) indicate the CME leading edge. The images are plotted using JHelioviewer (https://www.jhelioviewer.org/).
The panels (a) and (b) of Figure 7 depicts the changes in reconnection rate (RC flux rate) over time for both eruptions. We monitored the hottest region beneath the flux rope in the $r\textrm{-}\theta$ plane both before and after the eruption as shown in Figure 5). The peak reconnection rate (RC flux rate) for the first and second eruptions are $6.49\times 10^{17}$$\mathrm{Mx\,s^{-1}}$ and $4.10\times 10^{17}$$\mathrm{Mx\,s^{-1}}$, respectively. The panels (a) and (b) of Figure 7 indicate that the reconnection rate decreases after reaching its peak value, with the eruptions occurring near the maximum. This is similar to the observational data shown in Fig. 10b. After an eruption occurs, the reconnection sheet breaks down, making it challenging to track the field lines passing through the current sheet (CS).
III.2 Observed eruption
We now focus on finding signatures of such a correlation, if any, between reconnection flux and CME speed in solar observations. We study the temporal evolution of an M-class flare (SOL2011-SOL2011-08-04T03:41) which occurred in AR 11261 during 2011 August 4, and which had an associated CME. During this period, STEREO-A and STEREO-B were positioned at longitudinal separations of 101 and 92 (see inset of Figure 8), respectively, relative to the Sun-Earth line. This positioning allowed the two spacecrafts to provide near-limb views of Earth-directed CMEs. On the other hand, the location of AR 11261 was close to the solar disk center (between 30to 35 west) relative to Earth, offering reliable magnetic field data for the active region from the Heliospheric Magnetic Imager or HMI instrument (Schou et al., 2012). This provides a unique opportunity to study the co-temporal evolution of reconnection flux, estimated from on-disk SDO data, and the speed of the associated eruption, starting from the initiation height at lower corona as observed from near-limb observations by STEREO.
Figure 10: (a) Time profile of signed cumulative reconnection flux integrated over the positive (red) and negative (blue) magnetic polarities underlying the flare ribbons. (b) Time profile of unsigned instantaneous reconnection flux integrated over the positive (red solid circles) and negative (blue solid circles) magnetic polarities underlying the flare ribbons. The black solid line is the fitted profile over all the data points, including the red and blue solid circles. (c) Height time plot of the CME leading edge measured from STEREO EUVI, COR1 and COR2 for the same event as shown in Figure 8. (d) Velocity (in black) and acceleration (in blue) profiles of the CME leading edge derived from Panel (c).
We analyze the kinematics of the eruption associated with the M-class flare on 2011 August 04, by tracking the CME leading-edge height as observed by STEREO-A. We use three instruments from the Sun Earth Connection Coronal and Heliospheric Investigation (SECCHI; Howard et al. (2008)) onboard STEREO-A to track the early evolution of the Earth-directed CME eruption (see Figure 8). The instruments include the Extreme Ultraviolet Imager EUVI ($1.7R_{\odot}$), Wuelser et al. (2004) and the inner coronagraph COR1 ($1.3-4R_{\odot}$) and COR2 (Thompson et al., 2003).
We follow the method by (Kazachenko, 2023), to estimate the temporal profile of reconnecting flux. The observations of the flare ribbons are obtained from the Atmospheric Imaging Assembly (AIA; Lemen et al. (2012)) onboard SDO in 1600 Å channel. The associated full-disk magnetograms are obtained from the HMI instrument. During the recurrent eruptive M-class flare in AR 11261 on 2011 August 4, we identify the area swept out by the flare-ribbons observed in AIA 1600 Å images as shown in Figure 9. The half of the unsigned magnetic flux underlying the area associated with the brightening observed in AIA 1600 Å are recorded as the temporal profile of the re-connection flux. The kinematics of the CME eruption and the temporal variation in associated reconnection flux are shown in Figure 10.
III.3 Correlation between MFR acceleration and reconnection flux
@@
The panels (c) and (d) of Figure 7 show scatter plots of the MFR acceleration ($a_{r}$) versus the reconnection rate ($\phi_{RC}$) at different times during the evolution of the homologous synthetic eruptions. Before the eruption, the acceleration exhibits a monotonic relationship with the reconnection rate for the eruptive events. Notably, Pearson’s correlation coefficients before the eruption for the two events are 0.58 and 0.81, respectively, indicating a positive correlation between the reconnection rate and the acceleration. A similar qualitative relationship is also observed in the case of observational data (see panel (d) of Figure 10) where the acceleration of the erupting filament and the associated reconnection flux simultaneously increase, as indicated by the cyan shaded region. The correlation coefficients for this observed event before and after the eruption are 0.90 and 0.98, respectively. Furthermore, note that, in the lower corona, where the influence of solar wind is negligible, our results show that the speed of the flux rope decreases following the eruption, heralding the end of the acceleration phase. The deceleration of the CME is mainly because of adiabatic expansion and the resulting cooling of the magnetic cloud in a relatively large domain that we have used. This is illustrated in panels (e) & (f) of Figure 6 from our simulation model, which closely resembles panel (d) of Figure 10 based on our analysis of observational data. Similar findings have also been reported in Sarkar et al. (2019). Zhu et al. (2020) analyzed 60 CME-flare events and found correlation coefficients ranging from 0.68 to 0.99, indicating a strong correlation between acceleration and reconnection rate. Although our observational event is not included in their list, it also exhibits a similar correlation. Due to observational constraints (limited time cadence of STEREO EUVI), it was challenging to gather more data points during the initial acceleration phase of the eruption. The numerical analysis from our simulation complements the above-mentioned observational results, allowing us to determine the temporal evolution of the reconnection rate and acceleration and their correlation.
IV Summary and Discussion
@@
Our fully compressible MHD simulation using the Pencil Code presented here produces homologous CMEs due to the repeated formation and partial eruption of unstable flux ropes during the sustained emergence of highly twisted magnetic flux rope mimicking a newly formed coronal flux rope. Our results show that, with the continuous emergence of flux, a new current sheet (CS) forms above the same polarity inversion line (PIL) after the completion of a previous eruption, leading to a new eruption. This recurring formation and disruption of the coronal current sheet, driven by the ongoing emergence of flux, results in homologous eruptions.
@@
Our simulation began with the quasi-static emergence of a twisted flux rope at the lower boundary into a pre-existing coronal arcade. We found that the flux rope initially settles into a quasi-static rise phase, with an underlying sigmoid-shaped current layer developing beneath the flux rope. This current layer is likely the site for the formation of thin current sheets and magnetic reconnection (Titov, 2007; Aulanier et al., 2005; Savcheva et al., 2012). The reconnection in the current layer effectively adds twisted flux to the magnetic flux rope, allowing it to rise quasi-statically before undergoing a partial eruption. Subsequent flux emergence forms a new current channel, which exhibits similar behavior and erupts after the bodily emergence stops, or $v_{0}=0$.
We estimate by calculating the twist of the field line between the line tied ends that the first eruption remains stable to helical kink instability, while the second eruption is kink unstable. The subsequent CME, in our simulation, has a slightly lower speed and kinetic energy. Initially, the reconnection rate increases, followed by a gradual decrease, after each eruption. These trends are consistent with the observation of an eruptive event. Our simulation indicates that the reconnection process potentially influences the evolution of the magnetic flux rope. The reconnection rate shows a monotonic variation with the acceleration, especially before the eruption in our simulation. The importance of the present work lies in performing an MHD simulation to produce homologous CMEs and thereafter finding the correlation between the reconnection rate and CME acceleration as a function of time between pre to post-eruptive stage and further validating this relation using an observed event. We choose a event that appears in the database of both STEREO (as limb event) and SDO (as on-disk event) spacecrafts so as to enable us to estimate both speeds and reconnection flux as accurately as possible using these instruments. Despite the fact that the driving of the MHD simulation was not tailored to match the observed flare and the associated CME, a similar pattern has been detected in both cases between CME acceleration and reconnection rate, (see panel (b) of Figure 10), suggesting a consistent underlying process. The observed decrease in reconnection rate post-peak might be tied to the dynamic changes within the reconnection sheet during an eruption. As the reconnection sheet disintegrates, accurately measuring and tracking the magnetic field lines through the current sheet becomes challenging. This complexity impacts the calculated values of RC flux, reflecting the real-time physical changes occurring during an eruption. Moreover, these findings align with observational data, further validating the temporal relation between the reconnection rate and the onset of eruptions (panels b-d of Figure 10). This highlights the importance of considering reconnection rate during the rising phase of MFR evolution in studying the dynamics of solar eruptions and magnetic reconnection events. The correlation between RC flux rate and acceleration could therefore be a generic property of solar eruptions.
@@
Although our numerical simulation model does not include the solar wind, the inclusion of optically thin radiative cooling and explicit coronal heating makes it more reasonable than previously published homologous solar eruption models (Chatterjee & Fan, 2013; Bian et al., 2022). We account for localized heating due to the formation and dissipation of current sheets in the corona, as well as the redistribution of heat through field-aligned conduction. The lower boundary of our simulation domain is positioned at the coronal base rather than the photosphere, and therefore, our model does not simulate prominence formation.
@@
Although we employ an analytical configuration of the magnetic flux rope and include various physical processes in our model, this approach may not fully represent the realistic driving conditions of the solar eruptions. To address the complexity and evolution of the magnetic configuration in a real scenario, a detailed and accurate description of the evolving surface magnetic field is essential. Our future goal is to conduct detailed analysis of solar eruptions that are directly driven or constrained by photospheric magnetograms.
Acknowledgment